OpenStudy (ujjwal):

\(1+\sin x+ \sin^2 x +...... to \infty\)=4+3\(\sqrt3\), (\(0

OpenStudy (anonymous):

use trigonometric series a=1 q=sin x lql<1

OpenStudy (anonymous):

Geometric****

OpenStudy (ujjwal):

\[\frac{1}{1-\sin x}=4+3\sqrt3\]But none of the options seem to satisfy this.

OpenStudy (anonymous):

o.O

OpenStudy (ujjwal):

The closest to the answer seems to be \(\pi\)/3..

OpenStudy (anonymous):

that gives sin x=0.891